does the series ∑ 1/n⋅[1+(ln n)2] converge or diverge

n=1 below ∑

∞ on top of ∑






1
Expert's answer
2021-03-30T08:04:39-0400

"\sum _{n=1}^{\infty \:}\frac{1}{n}\left(1+\left(\ln \left(n\right)\right)^2\right)"

"As \space a_n=\frac{1}{n}\left(1+\left(\ln \left(n\right)\right)^2\right) ; a_x=\frac{1}{x}\left(1+\left(\ln \left(x\right)\right)^2\right)"

"L= \int_1^\infin \frac{1}{x}\left(1+\left(\ln \left(x\right)\right)^2\right)=\ln \left|x\right|+\frac{1}{3}\ln ^3\left(x\right)+C|_1^\infin =\infin"

Hence "\sum _{n=1}^{\infty \:}\frac{1}{n}\left(1+\left(\ln \left(n\right)\right)^2\right)" diverges